# A Survey of Agent Interoperability Protocols: Model Context Protocol (MCP), Agent Communication Protocol (ACP), Agent-to-Agent Protocol (A2A), and Agent Network Protocol (ANP)

### v1

@@ Abstract
Large language model (LLM)-powered autonomous agents demand robust, standardized protocols to integrate tools, share contextual data, and coordinate tasks across heterogeneous systems. Ad-hoc integrations are difficult to scale, secure, and generalize across domains. This survey examines four emerging agent communication protocols: Model Context Protocol (MCP), Agent Communication Protocol (ACP), Agent-to-Agent Protocol (A2A), and Agent Network Protocol (ANP), each addressing interoperability in distinct deployment contexts. MCP provides a JSON-RPC client-server interface for secure tool invocation and typed data exchange. ACP introduces REST-native messaging via multi-part messages and asynchronous streaming to support multimodal agent responses. A2A enables peer-to-peer task outsourcing through capability-based Agent Cards, facilitating enterprise-scale workflows. ANP supports open-network agent discovery and secure collaboration using decentralized identifiers (DIDs) and JSON-LD graphs. The protocols are compared across multiple dimensions, including interaction modes, discovery mechanisms, communication patterns, and security models. Based on the comparative analysis, a phased adoption roadmap is proposed: beginning with MCP for tool access, followed by ACP for multimodal messaging, A2A for collaborative task execution, and extending to ANP for decentralized agent marketplaces. This work provides a comprehensive foundation for designing secure, interoperable, and scalable ecosystems of LLM-powered agents.
Keywords Large Language Models (LLMs), Agent Communication, Interoperability Protocols, Model Context Protocol (MCP), Agent Communication Protocol (ACP), Agent-to-Agent Protocol (A2A), Agent Network Protocol (ANP), Autonomous Agents, Multimodal Messaging, Decentralized Identity (DID)
@@ 1 Introduction
  • Model Context Protocol (MCP): a JSON-RPC client–server interface for secure context ingestion and structured tool invocation [6, 7, 8].
  • Agent-to-Agent Protocol (A2A): a peer-to-peer framework using capability-based Agent Cards over HTTP and Server-Sent Events for enterprise-scale task orchestration [9].
  • Agent Communication Protocol (ACP): a REST-native performative messaging layer with multi-part messages, asynchronous streaming, and observability features for local multi-agent systems [10].
  • Agent Network Protocol (ANP): a decentralized discovery and collaboration protocol built on decentralized identifiers (DIDs) and JSON-LD graphs for open-internet agent marketplaces  [11, 12].
@@ 2 Challenges and Solutions in Agent Protocol Interoperability
Together, these protocols aim to transform fragmented AI ecosystems into robust, secure, and interoperable agent networks—scalable across organizational and vendor boundaries. See Table 7 for a detailed comparative overview.
@@ 3.5 Orchestration and Lightweight Agent Frameworks
CrewAI High-level “crew” abstractions for role assignment, subtask delegation, and message routing among agents [38]
@@ 5.5 Security Challenges and Mitigations Across the A2A Lifecycle
6 ACP Core Architecture
The Agent Communication Protocol (ACP) defines a layered, REST-native framework for interoperable AI agents. Its static architecture comprises distinct actors and protocol layers, each responsible for a well-defined set of functions.
@@ 6 ACP Core Architecture
6.1 ACP Architecture Overview
The Agent Communication Protocol (ACP) defines a streamlined three-role architecture designed to standardize discovery, invocation, and interaction among AI agents. This architecture ensures that clients can seamlessly locate agents, initiate structured task requests, and receive multimodal responses without requiring bespoke integrations.
At the entry point of this architecture is the Agent Client, which initiates communication by discovering agents through a published registry and composing structured requests in ACP compliant format. It encapsulates user intent into multi-part messages, manages session-level context if needed, and processes ordered response parts ranging from plain text to rich artifacts and binary data returned by agents.
At the center of the ACP system is the ACP Server, which acts as a protocol broker. It maintains the Agent Registry, a metadata catalog defined by the Agent Detail Schema, and enforces system-wide policies including authentication, authorization, and rate limiting. Upon receiving a client request, the ACP Server performs agent lookup and routing, ensures compliance with registered capabilities, and facilitates secure transport of response parts back to the client in the prescribed order.
The ACP Agent represents the execution endpoint, where domain-specific logic resides. It may operate as a stateless microservice or maintain session context to support multi-turn interactions. The agent ingests structured requests composed of ordered message parts each tagged with semantic metadata and processes them in accordance with the registered skill definitions. Upon task completion, the agent emits response parts that conform to ACP’s Message Structure specification, enabling a uniform and interpretable response pipeline.
Together, these three components establish a modular and interoperable framework for agent communication, promoting scalable deployment, loose coupling, and clear separation of discovery, orchestration, and execution logic.
@@ 7.4 Security Considerations Across the ANP Lifecycle
Strengths Tight LLM integration; resource injection Multimodal messaging; brokered registry; tool modularity Inter-agent negotiation; artifact-driven delegation DID-based trustless identity; AI-native protocol negotiation
@@ 9 Phased Adoption Roadmap for Agent Interoperability
This section presents a practical, multi-stage deployment strategy for agent interoperability based on protocol maturity, integration complexity, and domain-specific use cases. The roadmap helps organizations gradually adopt the most suitable agent communication standards while ensuring scalability, composability, and security.
@@ 9.1 Stage 1 – MCP for Tool Invocation
The initial phase involves adopting the Model Context Protocol (MCP) to enable secure and structured interaction between large language models (LLMs) and external tools or resources. MCP leverages a JSON-RPC-based client-server model, making it ideal for early agent integration where tool invocation, deterministic execution, and typed inputs/outputs are critical.
9.2 Stage 2 – ACP for Rich Interaction
Once foundational tool capabilities are in place, the Agent Communication Protocol (ACP) can be layered to support asynchronous, multimodal, and REST-native messaging. ACP introduces ordered message parts, flexible task schemas, and streaming support—features that enable richer agent conversations and integration with broader RESTful ecosystems.
@@ 9.3 Stage 3 – A2A for Enterprise Collaboration
In more complex enterprise environments, the Agent-to-Agent (A2A) protocol enables multi-agent workflows and task orchestration through structured capability cards and artifact exchanges. A2A supports dynamic discovery via Agent Cards and secure intra-organizational collaboration in trusted contexts.
@@ 10 Conclusion
As autonomous agents powered by large language models proliferate across domains, the demand for secure, modular, and interoperable communication grows increasingly urgent. This survey presented a structured analysis of four emerging protocols—MCP, ACP, A2A, and ANP—that each address distinct layers of agent interoperability. By unifying tool invocation, multimodal messaging, task coordination, and decentralized discovery, these protocols collectively form the foundation for scalable multi-agent systems. The comparative evaluation demonstrates that no single protocol suffices across all contexts; instead, a phased, complementary adoption strategy—beginning with MCP and progressing through ACP and A2A to ANP—offers a practical pathway for deploying agent ecosystems. Future research should explore protocol interoperability bridges, trust frameworks for agent collaboration, and standardized evaluation benchmarks to accelerate adoption and ensure resilience in real-world deployments. These foundational efforts will be critical for advancing the next generation of intelligent, networked agents.

### v2

@@ Abstract
Large language model (LLM)-powered autonomous agents demand robust, standardized protocols to integrate tools, share contextual data, and coordinate tasks across heterogeneous systems. Ad-hoc integrations are difficult to scale, secure, and generalize across domains. This survey examines four emerging agent communication protocols: Model Context Protocol (MCP), Agent Communication Protocol (ACP), Agent-to-Agent Protocol (A2A), and Agent Network Protocol (ANP), each addressing interoperability in distinct deployment contexts. MCP provides a JSON-RPC client-server interface for secure tool invocation and typed data exchange. ACP defines a general-purpose communication protocol over RESTful HTTP, supporting MIME-typed multipart messages and both synchronous and asynchronous interactions. Its lightweight and runtime-independent design enables scalable agent invocation, while features like session management, message routing, and integration with role-based and decentralized identifiers (DIDs).A2A enables peer-to-peer task delegation using capability-based Agent Cards, supporting secure and scalable collaboration across enterprise agent workflows. ANP supports open-network agent discovery and secure collaboration using W3C decentralized identifiers (DIDs) and JSON-LD graphs. The protocols are compared across multiple dimensions, including interaction modes, discovery mechanisms, communication patterns, and security models. Based on the comparative analysis, a phased adoption roadmap is proposed: beginning with MCP for tool access, followed by ACP for structured, multimodal messaging, session-aware interaction, and both online and offline agent discovery across scalable, HTTP-based deployments, A2A for collaborative task execution, and extending to ANP for decentralized agent marketplaces. This work provides a comprehensive foundation for designing secure, interoperable, and scalable ecosystems of LLM-powered agents.
Keywords Large Language Models (LLMs), Agent Communication, Interoperability Protocols, Model Context Protocol (MCP), Agent Communication Protocol (ACP), Agent-to-Agent Protocol (A2A), Agent Network Protocol (ANP), Autonomous Agents, Multimodal Messaging, Decentralized Identity (DID) , peer-to-peer
@@ 1 Introduction
  • Model Context Protocol (MCP): A JSON-RPC client–server interface for secure context ingestion and structured tool invocation. MCP streamline the integration of large language models (LLMs) with external data sources and tools. MCP addresses the challenges of fragmented and custom-built integrations by providing a universal, model-agnostic interface for AI systems to access and interact with diverse resources [6, 7, 8]. MCP was launched by Anthropic in Novemeber 2024.
  • Agent-to-Agent Protocol (A2A): Enable secure, structured, and interoperable collaboration between AI agents across platforms, vendors, and environments. A2A is designed to support peer-to-peer agent interactions using capability-based representations known as Agent Cards, which describe what an agent can do and how it can be securely invoked. A2A supports asynchronous, event-driven communication through HTTP and Server-Sent Events (SSE), making it suitable for distributed, scalable agent ecosystems, [9]. A2A was launched by Google in April 2025.
  • Agent Communication Protocol (ACP): ACP is a general-purpose protocol for agent communication that uses RESTful HTTP interfaces to support MIME-typed multipart messages and both synchronous and asynchronous interactions. This REST-based communication model enables lightweight, runtime-independent agent invocation, making ACP well-suited for scalable system integration. ACP includes structured session management, message routing, and a flexible authentication model that integrates with role-based access control (RBAC) and decentralized identity (DID) systems. It is compatible with a wide range of agent frameworks and deployment models, from lightweight stateless utilities to long-running, stateful services. ACP also supports agent discovery through runtime APIs, offline packaging, and manifest-based metadata. ACP was launched by IBM in March 2025 [10].
  • Agent Network Protocol (ANP): The Agent Network Protocol (ANP) is an open-source communication framework designed to enable secure, decentralized collaboration among AI agents across the open internet. Unlike traditional client-server architectures, ANP adopts a peer-to-peer (P2P) model, allowing agents to autonomously discover, authenticate, and interact with one another without centralized intermediaries  [11, 12].
@@ 2 Challenges and Solutions in Agent Protocol Interoperability
Together, these protocols aim to transform fragmented AI ecosystems into robust, secure, and interoperable agent networks scalable across organizational and vendor boundaries. See Table 7 for a detailed comparative overview.
@@ 3.5 Orchestration and Lightweight Agent Frameworks
CrewAI High-level crew abstractions for role assignment, subtask delegation, and message routing among agents [38]
@@ 5.5 Security Challenges and Mitigations Across the A2A Lifecycle
6 ACP Architecture
The Agent Communication Protocol (ACP) architecture is a modular, HTTP-based system designed to facilitate interaction between clients and agents. It is composed of three primary components: the ACP Client, the ACP Server, and one or more ACP Agents. Clients communicate with servers using standardized REST APIs, sending structured multi-modal messages that agents process and responds.
The ACP Client initiates communication by submitting requests in ACP-compliant format. It supports message composition using ordered message parts, session-based interactions for multi-turn workflows, and both synchronous and streaming execution modes. Error responses follow a unified structure, and standard HTTP authentication methods such as Bearer tokens, Basic Auth, and JWTs are supported.
The ACP Server acts as a middleware component, translating external HTTP requests into internal agent executions. It manages agent lifecycle phases including execution and optional session persistence. The server is stateless by default, making it compatible with load balancers and orchestration platforms such as Kubernetes. Secure communication is achieved via TLS, and observability is supported through OpenTelemetry-based tracing and metrics.
The ACP Agent is the execution component defined using a decorator-based configuration. Agents process structured requests consisting of ordered message parts and generate responses that conform to ACP’s message format. They support both stateless and session-aware operation, including features such as await/resume for interactive use cases. Agent metadata enables discovery through runtime APIs, offline packaging, or public manifest files.
Together, these three components form a framework that supports agent communication through standardized protocols, structured messaging, and flexible deployment strategies.
@@ 7.4 Security Considerations Across the ANP Lifecycle
Strengths Tight LLM integration; resource injection Multimodal messaging; brokered registry; tool modularity; Offline Agent Discovery Inter-agent negotiation; artifact-driven delegation DID-based trustless identity; AI-native protocol negotiation
@@ 9 Phased Adoption Roadmap for Agent Interoperability
This section presents a practical, multi-stage deployment strategy for agent interoperability based on protocol maturity, integration complexity, and use case alignment. The roadmap enables organizations to adopt suitable agent communication standards while supporting scalability, composability, and security.
@@ 9.1 Stage 1 – MCP for Tool Invocation
The initial phase involves adopting the Model Context Protocol (MCP) to enable structured and secure interaction between large language models (LLMs) and external tools or resources. MCP operates over a JSON-RPC-based client-server model and is well-suited for use cases focused on tool invocation, deterministic execution, and typed input/output. This stage establishes a foundation for context enrichment in single-model systems.
9.2 Stage 2 – ACP for Agent Communication
The second phase introduces the Agent Communication Protocol (ACP), a general-purpose, REST-native protocol designed for asynchronous and synchronous communication between independent agents. ACP supports MIME-typed multipart messages, ordered message parts, and streaming. These capabilities enable standardized messaging across agents built on different frameworks, facilitating collaboration and interoperability across organizational and technological boundaries.
@@ 9.3 Stage 3 – A2A for Enterprise Collaboration
In enterprise-oriented deployments, the Agent-to-Agent (A2A) protocol facilitates multi-agent interaction through Agent Cards and artifact exchanges. It supports dynamic discovery and structured messaging using predefined types, making it well-suited for coordinated workflows among stateless or stateful agents within trusted organizational contexts.
@@ 10 Conclusion
As autonomous agents powered by large language models proliferate across domains, the demand for secure, modular, and interoperable communication grows increasingly urgent. This survey presented a structured analysis of four emerging protocols MCP, ACP, A2A, and ANP that each address distinct layers of agent interoperability. By unifying tool invocation, multimodal messaging, task coordination, and decentralized discovery, these protocols collectively form the foundation for scalable multi-agent systems. The comparative evaluation demonstrates that no single protocol suffices across all contexts; instead, a phased, complementary adoption strategy—beginning with MCP and progressing through ACP and A2A to ANP offers a practical pathway for deploying agent ecosystems. Future research should explore protocol interoperability bridges, trust frameworks for agent collaboration, and standardized evaluation benchmarks to accelerate adoption and ensure resilience in real-world deployments. These foundational efforts will be critical for advancing the next generation of intelligent, networked agents.